# DRAGON: Dynamic RAG Benchmark On News

### v1

@@ 1 Introduction
(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
@@ 3 System Demo
Architecture of the Dynamic Benchmark
@@
Data Acquisition and Processing: We have maintained a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx. A for the news sources list) on a daily basis. Newly parsed content is synchronized into our storage. To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
@@
QA Dataset Formation: The process of creating the questions and answers based on the updated increment of news data is described in detail in Sec. 4. The pipeline transforms generated QA pairs into several HF datasets, which form the core of our benchmark:
@@ 4 Methodology
The Data Generation pipeline (Fig. 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation. The KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph. The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
@@
All the extracted entities are used to match with entities from Wikidata555We use Russian subgraph of Wikidata as a base one. Vrandečić and Krötzsch (2014). To perform this, we search by API in this open knowledge base for the entities extracted in the previous step. Afterward, the found entities are vectorized using an embedding model and saved as a vector database. Then, for each entity extracted in the first step, we list five of the closest entity candidates in the vector representation base.
@@ 6 Experiments
Retrieval evaluation results presented in Tab. 2 demonstrate consistently strong performance across all evaluated retriever models. Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par except for MRR, where E5 Mistral7b Instruct beats its competitor. However, mE5lInstruct and FRIDA also perform competitively.
@@ Appendix I Human Evaluation Interface
Appendix J LLM-as-Judge RAG Evaluation
This section provides a detailed description of the LLM-as-Judge criteria used to evaluate RAG systems. To build a comprehensive and interpretable set of metrics, Evaluation Targets provided by Yu et al. (2024) are utilized. For each Evaluation Target, we select several criteria from the POLLUX Martynov et al. (2025) set of criteria:

### v2

@@ 1 Introduction
(iii) We launch a public, regularly updated leaderboard333https://huggingface.co/spaces/ai-forever/rag-leaderboard for recurrent evaluation to support reproducible and community-driven research.
@@ 3 System Demo
Dynamic Benchmark Architecture
@@
Data Acquisition and Processing: We maintain a dedicated set of parsers which periodically crawl a selection of news sources recognized as popular news websites in Russia (see Appx. A for the news sources list) on a daily basis. Newly parsed content is synchronized into our storage. To avoid redundancy and ensure incremental updates, a scheduled automated job identifies differences with the previous dataset revision and extracts updated segments for downstream processing.
@@
QA Dataset Formation: The process of creating the questions and answers based on the updated increment of news data is described in detail in Sec. 4. The pipeline transforms the generated QA pairs into several HF datasets, which form the core of our benchmark:
@@ 4 Methodology
The Data Generation pipeline (Fig. 3) consists of 2 main stages preceded by preliminary data preprocessing: KG Extraction and Question Generation. KG Extraction retrieves factual information from texts and preserves the most specific and fresh facts in the form of a Knowledge Graph. The Question Generation module samples subgraphs of a certain structure to generate a question-answer pair with LLM.
@@
All the extracted entities are used to match with entities from Wikidata555We use Russian subgraph of Wikidata as a base one. Vrandečić and Krötzsch (2014). To do this, we search by API in this open knowledge base for the entities extracted in the previous step. Afterward, the found entities are vectorized using an embedding model and saved as a vector database. Then, for each entity extracted in the first step, we list five of the closest entity candidates in the vector representation base.
@@ 6 Experiments
Retrieval evaluation results presented in Tab. 2 demonstrate consistently strong performance across all evaluated retriever models. Among them, Qwen3Embedding 8B and E5 Mistral7b Instruct achieve the highest scores, performing practically on par. mE5lInstruct also performs competitively. As for FRIDA, it also demonstrates strong performance but its results are slightly inferior to those of the competitors.
@@ Appendix I Human Evaluation Interface
Appendix J LLM-as-Judge RAG Evaluation Criteria
This section provides a detailed description of the LLM-as-Judge criteria used to evaluate RAG systems. To build a comprehensive and interpretable set of metrics, Evaluation Targets provided by Yu et al. (2024) are utilized. For each Evaluation Target, we select several criteria from the POLLUX set of criteria: